# Abelian surfaces over finite fields containing no curves of genus $3$ or less

### v3

@@ 5. Polarisations of degree $4$ on abelian surfaces
This implies that $[\mathcal{O}/\mathfrak{M}]_{R}=[R/\mathfrak{l}R]_{R}$. If we denote the image of $[\mathcal{O}/\mathfrak{M}]_{\mathcal{O}}$ in $H(\mathcal{O})$ by $y_{\mathfrak{M}}$, we get $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. By Proposition 3.5, since $\mathcal{O}/\mathfrak{M}\simeq\mathbb{F}_{4}$, we get that $\mathfrak{m}$ is inert in $K$. This means that the Artin symbol $\left(\frac{K/K^{+}}{\mathfrak{m}}\right)$ is a generator of $\mathop{\rm Gal}\nolimits(K/K^{+})$. We now claim that $K/K^{+}$ is unramified. If not, then by Lemma 3.1 and Lemma 3.2, we would get that there is unique maximal ideal $\mathfrak{m}$ of $\mathcal{O}^{+}$ above $2$ which ramifies in $\mathcal{O}$, say $\mathfrak{m}\mathcal{O}=\mathfrak{M}^{2}$. For such $\mathfrak{M}$ we would have $\mathcal{O}/\mathfrak{M}\simeq\mathbb{F}_{2}$. Since we know that every maximal ideal of $\mathcal{O}$ above $\mathfrak{l}$ has residue field isomorphic to $\mathbb{F}_{4}$, we deduce that $K/K^{+}$ is unramified. Then, by Proposition 4.8, we get that $\chi(y_{\mathfrak{M}})\neq 0$ in $\mathcal{B}(\mathcal{O})$. So $(x_{\mathfrak{l}},0)$ is not in the image of $(N,-\chi)$ and, therefore, there is no $R$-module of order $4$ which is attainable. ∎
In Theorem 5.7, we wrap-up the discussion above. Moreover, in certain cases with $\mathcal{C}$ ordinary, we show that if there is an abelian variety with polarisation with kernel of order $4$ then there is an abelian variety with polarisation with kernel of order $4$ and maximal endomorphism ring. To do so, we use results from [13]. The construction of $\mathcal{B}(\mathcal{O})$ in [13] is slightly different form the one in [14] that we have presented above, but certainly equivalent in view of Proposition 4.8 and [13, Prop. 10.1].
@@
Clearly strongly attainable implies attainable, while the contrary does not hold in general. A characterisation of strongly attainable $R$-modules is given in [13, Prop. 5.7], repeated below as Proposition 5.5 for convenience. Such characterisation relates modules with a particular element $I_{K,\Phi}$ of $\mathcal{B}(\mathcal{O})$ which depends on a choice of a CM-type $\Phi$ of $K$. The definition of $I_{K,\Phi}$ is given in [13, Def. 5.2]. For our purposes, it sufficient to know whether it is trivial or not in $\mathcal{B}(\mathcal{O})$.
@@
Assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ is ordinary. Then $\mathcal{C}$ does not contain a principally polarised abelian variety. The result follows from [13, Prop. 11.3] and [13, Cor. 11.4].
Assume from now that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ is ordinary. If $K/K^{+}$ is ramified then the $I_{K,\Phi}=0$ by [13, Prop. 11.1]. If $K/K^{+}$ is unramified and there is no maximal ideal of $\mathcal{O}^{+}$ dividing $\pi-q/\pi$ which stays inert in $K/K^{+}$ then the result follows, again, by combining [13, Prop. 11.3] and [13, Cor. 11.4] and the observation that $\mathcal{C}$ contains a principally polarised abelian variety. ∎
@@
[13, Prop. 5.7] Let $M$ be finite $R$-module. Then $M$ is strongly attainable if and only if $M\simeq\mathcal{O}/\mathfrak{a}\mathcal{O}$ for some $\mathcal{O}^{+}$-ideal $\mathfrak{a}\subseteq\mathcal{O}^{+}$ such that the image in $\mathcal{B}(\mathcal{O})$ of the class of $\mathcal{O}^{+}/\mathfrak{a}$ in $G(\mathop{\rm Mod}\nolimits_{\mathcal{O}^{+}})$ equals $I_{K,\Phi}$.
@@
      $b=1-2q$ and $q$ is odd, if $\mathcal{C}$ is ordinary; $q$ is even, if $\mathcal{C}$ is non-ordinary.
@@
The statement Theorem 5.7.ii can be deduced from the claim contained in [12, page 155, 2nd par. section 3], stated without a detailed proof.
6. Computing isomorphism classes admitting a polarisation of degree $4$
@@
Using the aforementioned algorithm, we count the number of polarisations of degree $4$, for every isogeny class in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ for a fixed range of $q$. For the full output, see https://raw.githubusercontent.com/stmar89/Genus3Data/main/table_output.txt. We remark that in the maximal endomorphism ring case, the ratio of isomorphism classes of abelian surfaces admitting a polarisation of degree $4$ is always 00 or a power of $1/2$.
@@ 7. Genus $3$ curves lying on abelian surfaces
Here we mainly follow [5, Sec. 1.2], with some minor modifications since we are working over finite fields instead of over $\mathbb{C}$. By Theorem 2.2, the curve $C$ induces a polarisation $\lambda_{\mathcal{L}}:A\to A^{\vee}$ of degree $4$, where $\mathcal{L}=\mathcal{L}(C)$ is the ample line bundle defined by $C$. Let $\mathcal{B}$ be the base locus of $\mathcal{L}$, that is, =$\mathcal{B}=\{P\in A\,\mid\,s(P)=0\text{ for any section }s\in H^{0}(A,\mathcal{L})\}$ [10, II, Lem. 7.8]. Fix some $P_{0}\in\mathcal{B}$. Since $\ker\lambda_{\mathcal{L}}=\{P\in A\,\mid\,t_{P}^{*}\mathcal{L}\simeq\mathcal{L}\}$ [21, Ch. I, Sec. 8] we have, for any and $P\in\ker\lambda_{\mathcal{L}}$, $s(P_{0}+P)=t_{P}^{*}s(P_{0})=s(P_{0})=0$. Thus $P_{0}+\ker\lambda_{\mathcal{L}}\subseteq\mathcal{B}$. Since we are in characteristic different from $2$, $\ker\lambda_{\mathcal{L}}$ has order equal to $4$, the degree of $\lambda_{\mathcal{L}}$. On the other hand, $\mathcal{B}$ contains at most $C.C=4$ points. Hence, $P_{0}+\ker\lambda_{\mathcal{L}}$ and $\mathcal{B}$ coincide as sets of geometric points. Consider now the involution $\iota:A\to A$ given by $P\mapsto-P+2P_{0}$. Since the points in $\ker\lambda_{\mathcal{L}}$ are $2$-torsion points, the points in $\mathcal{B}$ are fixed points of $\iota$. Furthermore, we have $\iota^{*}\mathcal{L}\simeq\mathcal{L}$, and in particular $\iota C=C$ [5, Prop. 1.6]. Then one can lift $\iota$ to $C$ as an involution, and since $C$ is smooth, $\iota$ has exactly four fixed points on $C$, namely the points in $\mathcal{B}$ [5, 1.7]. Then, by the genus formula, the quotient $E=C/\iota$ is an elliptic curve, that is, $C$ is a double cover of $E$. Note that $\iota$ is not necessarily an $\mathbb{F}_{q}$-rational map: this will depend on whether $\mathcal{B}$ contains a rational point or not, that is, whether one can take $P_{0}$ to be rational. However, the fact that $\mathcal{B}$ is rational as a set implies that it is a (necessarily disjoint) union of orbits under the action of the Frobenius endomorphism. A point is of degree $n$ if the size of its orbit is $n$. As the size of $\mathcal{B}$ is $4$, the degree of its points is at most $4$. Hence the point $P_{0}$ defining the involution $\iota$ will have degree at most $4$. So $\iota$ and $E$ are defined over an extension of $\mathbb{F}_{q}$ of degree at most $4$. ∎
For the rest of the section, we fix an abelian surface $A$ defined over $\mathbb{F}_{q}$ with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$, that is, which is simple and not isogenous to a Jacobian, or equivalently, not containing absolutely irreducible curves of geometric genus 00, $1$ or $2$ by Theorem 1.4. We suppose that there is an absolutely irreducible smooth genus $3$ curve $C$ defined over $\mathbb{F}_{q}$ that lies on $A$, and we want to study that curve.
@@
Finally, we present bounds on the number of rational points that an absolutely irreducible smooth genus $3$ curve lying on an abelian surface with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$ can have. To start with, we recall and extend a result from [9] on the number of rational points on curves over abelian surfaces.
@@
Let $A$ be an abelian surface defined over $\mathbb{F}_{q}$ with trace $-a$, that is, with Weil polynomial $t^{4}+at^{3}+bt^{2}+qat+q^{2}$. Let $C$ be an absolutely irreducible curve defined over $\mathbb{F}_{q}$, of arithmetic genus $p_{a}(C)=p_{a}$, lying on $A$. Then
|#$|\#C(\mathbb{F}_{q})-(q+1+a)|\leq|p_{a}-2|\lfloor 2\sqrt{q}\rfloor.$
@@
The upper bound on #$\#C(\mathbb{F}_{q})$ is [9, Thm. 4]. The proof can be adapted to obtain also the lower bound on #$\#C(\mathbb{F}_{q})$. In what follows, we borrow the notation from the proof of [9, Thm. 4].
If $p_{a}=1$, then we know that the curve $C$ is elliptic, of trace say $-e$. Write $a=e+x_{2}$ for some integer $-\lfloor 2\sqrt{q}\rfloor\leq x_{2}\leq\lfloor 2\sqrt{q}\rfloor$. Hence,
#$\#C(\mathbb{F}_{q})=q+1+e=q+1+a-x_{2}\geq q+1+a-\lfloor 2\sqrt{q}\rfloor.$
Suppose now $p_{a}\geq 2$. By [3, Prop. 2.3], we know that #$\#C(\mathbb{F}_{q})-\#\tilde{C}(\mathbb{F}_{q})\geq g-p_{a}$, where $\tilde{C}$ is the normalisation of $C$ and $g$ is its genus. Following the reasoning in the proof of [9, Thm. 4, Eq. (5)], there exist real numbers $x_{3},\dots,x_{g}$ such that
#$\#\tilde{C}(\mathbb{F}_{q})=q+1+a+\sum_{i=3}^{g}x_{i}\,\text{ and }\,\sum_{i=3}^{g}x_{i}\geq-(g-2)\lfloor 2\sqrt{q}\rfloor.$
@@
In particular, we see that an absolutely irreducible curve of genus $3$ lying on a simple abelian surface not isogenous to the Jacobian of an absolutely irreducible smooth genus $2$ curve is far from reaching the Serre–Weil bound which states |#$|\#C(\mathbb{F}_{q})-(q+1)|\leq 3\lfloor 2\sqrt{q}\rfloor$.
@@ Acknowledgments
The first author is supported by the grants ANR-21-CE39-0009-BARRACUDA and ANR-22-CPJ2-0047-01. The third author is supported by Nederlandse Organisatie voor Wetenschappelijk Onderzoek, grant number VI.Veni.202.107, by Agence Nationale de la Recherche under the MELODIA project, grant number ANR-20-CE40-0013, and by Marie-Skłodowska-Curie Actions - Postdoctoral Fellowships 2023 (project 101149209 - AbVarFq).

### v4

@@ 5. Polarisations of degree $4$ on abelian surfaces
This implies that $\mathcal{O}/\mathfrak{M}]_{R}=[R/\mathfrak{l}R]_{R}$. If we denote the image of $\mathcal{O}/\mathfrak{M}]_{\mathcal{O}}$ in $H(\mathcal{O})$ by $y_{\mathfrak{M}}$, we get $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. By Proposition 3.5, since $\mathcal{O}/\mathfrak{M}\simeq\mathbb{F}_{4}$, we get that $\mathfrak{m}$ is inert in $K$. This means that the Artin symbol $\left(\frac{K/K^{+}}{\mathfrak{m}}\right)$ is a generator of $\mathop{\rm Gal}\nolimits(K/K^{+})$. We now claim that $K/K^{+}$ is unramified. If not, then by Lemma 3.1 and Lemma 3.2, we would get that there is unique maximal ideal $\mathfrak{m}$ of $\mathcal{O}^{+}$ above $2$ which ramifies in $\mathcal{O}$, say $\mathfrak{m}\mathcal{O}=\mathfrak{M}^{2}$. For such $\mathfrak{M}$ we would have $\mathcal{O}/\mathfrak{M}\simeq\mathbb{F}_{2}$. Since we know that every maximal ideal of $\mathcal{O}$ above $\mathfrak{l}$ has residue field isomorphic to $\mathbb{F}_{4}$, we deduce that $K/K^{+}$ is unramified. Then, by Proposition 4.8, we get that $\chi(y_{\mathfrak{M}})\neq 0$ in $\mathcal{B}(\mathcal{O})$. So $(x_{\mathfrak{l}},0)$ is not in the image of $(N,-\chi)$ and, therefore, there is no $R$-module of order $4$ which is attainable. ∎
In Theorem 5.7, we wrap-up the discussion above. Moreover, in certain cases with $\mathcal{C}$ ordinary, we show that if there is an abelian variety with polarisation with kernel of order $4$ then there is an abelian variety with polarisation with kernel of order $4$ and maximal endomorphism ring. To do so, we use results from [15]. The construction of $\mathcal{B}(\mathcal{O})$ in [15] is slightly different form the one in [16] that we have presented above, but certainly equivalent in view of Proposition 4.8 and [15, Prop. 10.1].
@@
Clearly strongly attainable implies attainable, while the contrary does not hold in general. A characterisation of strongly attainable $R$-modules is given in [15, Prop. 5.7], repeated below as Proposition 5.5 for convenience. Such characterisation relates modules with a particular element $I_{K,\Phi}$ of $\mathcal{B}(\mathcal{O})$ which depends on a choice of a CM-type $\Phi$ of $K$. The definition of $I_{K,\Phi}$ is given in [15, Def. 5.2]. For our purposes, it sufficient to know whether it is trivial or not in $\mathcal{B}(\mathcal{O})$.
@@
Assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ is ordinary. Then $\mathcal{C}$ does not contain a principally polarised abelian variety. The result follows from [15, Prop. 11.3] and [15, Cor. 11.4].
Assume from now that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ is ordinary. If $K/K^{+}$ is ramified then the $I_{K,\Phi}=0$ by [15, Prop. 11.1]. If $K/K^{+}$ is unramified and there is no maximal ideal of $\mathcal{O}^{+}$ dividing $\pi-q/\pi$ which stays inert in $K/K^{+}$ then the result follows, again, by combining [15, Prop. 11.3] and [15, Cor. 11.4] and the observation that $\mathcal{C}$ contains a principally polarised abelian variety. ∎
@@
[15, Prop. 5.7] Let $M$ be finite $R$-module. Then $M$ is strongly attainable if and only if $M\simeq\mathcal{O}/\mathfrak{a}\mathcal{O}$ for some $\mathcal{O}^{+}$-ideal $\mathfrak{a}\subseteq\mathcal{O}^{+}$ such that the image in $\mathcal{B}(\mathcal{O})$ of the class of $\mathcal{O}^{+}/\mathfrak{a}$ in $G(\mathop{\rm Mod}\nolimits_{\mathcal{O}^{+}})$ equals $I_{K,\Phi}$.
@@
      b=12qb=1-2q and $q$ is odd, if $\mathcal{C}$ is ordinary; $q$ is even, if $\mathcal{C}$ is non-ordinary.
@@
The statement Theorem 5.7.ii can be deduced from the claim contained in [14, page 155, 2nd par. section 3], stated without a detailed proof.
6. Computing isomorphism classes admitting a polarisation of degree 44
@@
Using the aforementioned algorithm, we count the number of polarisations of degree $4$, for every isogeny class in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ for a fixed range of $q$. For the full output, see https://raw.githubusercontent.com/stmar89/Genus3Data/main/table_output.txt. We remark that in the maximal endomorphism ring case, the ratio of isomorphism classes of abelian surfaces admitting a polarisation of degree $4$ is always 0 or a power of $1/2$.
@@ 7. Genus $3$ curves lying on abelian surfaces
Here we mainly follow [5, Sec. 1.2], with some minor modifications since we are working over finite fields instead of over $\mathbb{C}$. By Theorem 2.2, the curve $C$ induces a polarisation $\lambda_{\mathcal{L}}:A\to A^{\vee}$ of degree $4$, where $\mathcal{L}=\mathcal{L}(C)$ is the ample line bundle defined by $C$. Let $\mathcal{B}$ be the base locus of $\mathcal{L}$, that is, $\mathcal{B}=\{P\in A\,\mid\,s(P)=0\text{ for any section }s\in H^{0}(A,\mathcal{L})\}$ [10, II, Lem. 7.8]. Fix some $P_{0}\in\mathcal{B}$. Since $\ker\lambda_{\mathcal{L}}=\{P\in A\,\mid\,t_{P}^{*}\mathcal{L}\simeq\mathcal{L}\}$ [21, Ch. I, Sec. 8] we have, for any and $P\in\ker\lambda_{\mathcal{L}}$, $s(P_{0}+P)=t_{P}^{*}s(P_{0})=s(P_{0})=0$. Thus $P_{0}+\ker\lambda_{\mathcal{L}}\subseteq\mathcal{B}$. Since we are in characteristic different from $2$, $\ker\lambda_{\mathcal{L}}$ has order equal to $4$, the degree of $\lambda_{\mathcal{L}}$. On the other hand, $\mathcal{B}$ contains at most $C.C=4$ points. Hence, $P_{0}+\ker\lambda_{\mathcal{L}}$ and $\mathcal{B}$ coincide as sets of geometric points. Consider now the involution $\iota:A\to A$ given by $P\mapsto-P+2P_{0}$. Since the points in $\ker\lambda_{\mathcal{L}}$ are $2$-torsion points, the points in $\mathcal{B}$ are fixed points of $\iota$. Furthermore, we have $\iota^{*}\mathcal{L}\simeq\mathcal{L}$, and in particular $\iota C=C$ [5, Prop. 1.6]. Then one can lift $\iota$ to $C$ as an involution, and since $C$ is smooth, $\iota$ has exactly four fixed points on $C$, namely the points in $\mathcal{B}$ [5, 1.7]. Then, by the genus formula, the quotient $E=C/\iota$ is an elliptic curve, that is, $C$ is a double cover of $E$. Note that $\iota$ is not necessarily an $\mathbb{F}_{q}$-rational map: this will depend on whether $\mathcal{B}$ contains a rational point or not, that is, whether one can take $P_{0}$ to be rational. However, the fact that $\mathcal{B}$ is rational as a set implies that it is a (necessarily disjoint) union of orbits under the action of the Frobenius endomorphism. A point is of degree $n$ if the size of its orbit is $n$. As the size of $\mathcal{B}$ is $4$, the degree of its points is at most $4$. Hence the point $P_{0}$ defining the involution $\iota$ will have degree at most $4$. So $\iota$ and $E$ are defined over an extension of $\mathbb{F}_{q}$ of degree at most $4$. ∎
For the rest of the section, we fix an abelian surface $A$ defined over $\mathbb{F}_{q}$ with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$, that is, which is simple and not isogenous to a Jacobian, or equivalently, not containing absolutely irreducible curves of geometric genus 0, 11 or $2$ by Theorem 1.4. We suppose that there is an absolutely irreducible smooth genus $3$ curve $C$ defined over $\mathbb{F}_{q}$ that lies on $A$, and we want to study that curve.
@@
Finally, we highlight that if an absolutely irreducible genus $3$ curve can be embedded in an abelian surface with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$, then its number of rational points is far from the Serre-Weil bound. This illustrates how embedding entails arithmetic constraints on curves. To start with, we recall and extend a result from [9] on the number of rational points on curves over abelian surfaces.
@@
Let $A$ be an abelian surface defined over $\mathbb{F}_{q}$ with trace a-a, that is, with Weil polynomial $t^{4}+at^{3}+bt^{2}+qat+q^{2}$. Let $C$ be an absolutely irreducible curve defined over $\mathbb{F}_{q}$, of arithmetic genus $p_{a}(C)=p_{a}$, lying on $A$. Then
$|\#C(\mathbb{F}_{q})-(q+1+a)|\leq|p_{a}-2|\lfloor 2\sqrt{q}\rfloor.$
@@
The upper bound on $C(\mathbb{F}_{q})$ is [9, Thm. 4]. The proof can be adapted to obtain also the lower bound on $C(\mathbb{F}_{q})$. In what follows, we borrow the notation from the proof of [9, Thm. 4].
If $p_{a}=1$, then we know that the curve $C$ is elliptic, of trace say e-e. Write $a=e+x_{2}$ for some integer $\lfloor 2\sqrt{q}\rfloor\leq x_{2}\leq\lfloor 2\sqrt{q}\rfloor$. Hence,
$C(\mathbb{F}_{q})=q+1+e=q+1+a-x_{2}\geq q+1+a-\lfloor 2\sqrt{q}\rfloor.$
Suppose now $p_{a}\geq 2$. By [3, Prop. 2.3], we know that $C(\mathbb{F}_{q})-\#\tilde{C}(\mathbb{F}_{q})\geq g-p_{a}$, where $\tilde{C}$ is the normalisation of $C$ and $g$ is its genus. Following the reasoning in the proof of [9, Thm. 4, Eq. (5)], there exist real numbers $x_{3},\dots,x_{g}$ such that
$\tilde{C}(\mathbb{F}_{q})=q+1+a+\sum_{i=3}^{g}x_{i}\,\text{ and }\,\sum_{i=3}^{g}x_{i}\geq-(g-2)\lfloor 2\sqrt{q}\rfloor.$
@@
In particular, we see that an absolutely irreducible curve of genus $3$ lying on a simple abelian surface not isogenous to the Jacobian of an absolutely irreducible smooth genus $2$ curve is far from reaching the Serre–Weil bound which states $|\#C(\mathbb{F}_{q})-(q+1)|\leq 3\lfloor 2\sqrt{q}\rfloor$.
@@ Acknowledgments
The first author is supported by the grants ANR-21-CE39-0009-BARRACUDA, ANR-22-CPJ2-0047-01, and PEPS “Jeunes chercheurs et jeunes chercheuses” 2025 from Insmi. The third author is supported by Nederlandse Organisatie voor Wetenschappelijk Onderzoek, grant number VI.Veni.202.107, by Agence Nationale de la Recherche under the MELODIA project, grant number ANR-20-CE40-0013, and by Marie-Skłodowska-Curie Actions - Postdoctoral Fellowships 2023 (project 101149209 - AbVarFq).